# On the Fundamental Impossibility of Hallucination Control in Large Language Models

### v5

@@ 4.1.1 Limits of Knowledge and Reasoning
For any $A\subseteq\mathcal{K}$ and context $Q\subseteq\mathcal{K}$ there exists
$\mu_{\infty}(A|Q)=\lim_{C\to\infty}\mu_{C}(A|Q).$ (5)
@@
Since $\mu_{C}(A|Q)\leq\mu_{D}(A|Q)$ when $C<D$ and $\mu_{C}$ is bounded on finite $\mathcal{K}$, the monotonic increasing and bounded sequence $\{\mu_{C}(A|Q)\}_{C}$ converges with $C\to\infty$ for any subset of knowledge space. ∎
It may be surprising to discover that the existence of that limit does not prohibit emergent information. We can learn and obtain useful information without violating that fundamental limitation above. To explain that critical and controversial property of information processing, we first introduce the emergence operator $\mathcal{E}_{C}$ to describe knowledge becoming accessible through (computationally) bounded reasoning. Then we can study the implications in more details.
@@
Let $\mathcal{F}_{C}$ be a complete set of all computational transformations with cost (complexity) lower than or equal to $C$, such that $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ when $C<D$ and such that for any $f,g\in\mathcal{F}_{\infty}=\bigcup_{C}\mathcal{F}_{C}$ there exists $h\in\mathcal{F}_{\infty}$ for which $g(A\cup f(A))=h(A)$. Also, we call $f$ monotone when it is monotone with respect to set inclusion.
The emergence operator is a mapping $\mathcal{E}_{C}\colon\mathcal{P}(\mathcal{K})\times\mathcal{P}(\mathcal{K})\rightarrow\mathcal{P}(\mathcal{K})$ defined as follows:
$\mathcal{E}_{C}(A|Q)=A\cup\{k\in\mathcal{K}\mid k=f(A)\in\mathcal{R}_{C}(Q)\ \text{for some monotone}\ f\in\mathcal{F}_{C}\}.$ (6)
Furthermore, to represent reasoning with unlimited computational budget, we define:
$\mathcal{E}_{\infty}(A)=A\cup\{k\in\mathcal{K}\mid k=f(A)\in\mathcal{R}_{C}(Q)\ \text{for some monotone}\ f\in\mathcal{F}_{\infty}\}.$ (7)
Finally, for $n\geq 0$ we have:
@@
Philosophically, the emergence operator describes reasoning as explicitation, i.e., as the process of making latent implications manifest. When $A\subseteq\mathcal{E}_{C}(A|Q)$, the discovered additional elements represent knowledge that was implicit in $A$ but required computational work $C$ to become accessible. That aligns with the pragmatist or constructivist view that understanding involves active construction rather than passive reception.
When $\mathcal{E}_{C}(A|Q)=A\cup B$, we say that $B$ emerges from reasoning based on $A$. Then $A$ and $B$ are computationally dependent and we write $A\not\perp_{C}B$. In such a case, it follows from the properties of $\mu_{C}$ that we can gain useful information. Whenever $A\subseteq A\cup B$, we have
@@
The emergence operator is monotonic, which means the reasoning process it describes is also monotonic. Whenever $A\subseteq B$, we must have $\mathcal{E}_{C}(A|Q)\subseteq\mathcal{E}_{C}(B|Q)$. Also, we preserve insight monotonicity, because by definition with $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ for $C<D$, we have $\mathcal{E}_{C}(A|Q)\subseteq\mathcal{E}_{D}(A|Q)$. By design, it also guarantees existence of a fixed point to which the composition of reasoning rules converges.
@@
For any $A\subseteq\mathcal{K}$ and budget $C>0$, there exists a fixed point of the emergence operator $\mathcal{E}_{C}$:
$A^{*}_{C}=\bigcup_{n\geq 0}^{|\mathcal{K}|}\mathcal{E}_{C}^{(n)}(A|Q)=\mathcal{E}_{C}(A^{*}_{C}|Q).$ (10)
@@
Since $(\mathcal{P}(\mathcal{K}),\subseteq)$ is a complete lattice and $\mathcal{E}_{C}$ is monotone, by the Knaster-Tarski fixed point theorem Tarski1955ALF there exists a fixed point $A^{*}_{C}$ of reasoning with $\mathcal{E}_{C}$. Finiteness of $\mathcal{K}$ ensures convergence within $|\mathcal{K}|$ steps. ∎
Let us apply the results above to show how $\mu_{\infty}$ can be constructed with $\mathcal{E}_{\infty}$. One example exploits the reasoning operations as follows:
$\mu_{\infty}(A|Q)=\log|\mathcal{K}|\cdot|\bigcup_{n=0}^{|\mathcal{K}|}\mathcal{E}^{(n)}_{\infty}(A|Q)|.$ (11)
That $\mu_{\infty}$ satisfies axioms demanded in Definition 5. Clearly, $\mu_{\infty}$ is bounded and
$\mu_{\infty}(\emptyset|Q)=\log|\mathcal{K}|\cdot|\bigcup_{n=0}^{|\mathcal{K}|}\mathcal{E}^{(n)}_{\infty}(\emptyset|Q)|=\log|\mathcal{K}|\cdot|\emptyset|=0.$
If $A\subseteq B$, then $\mathcal{E}_{\infty}(A|Q)\subseteq\mathcal{E}_{\infty}(B|Q)$ implies $\mu_{\infty}(A|Q)\leq\mu_{\infty}(B|Q)$. When $A\perp_{\infty}B$, no element of $A^{*}_{\infty}$ can be derived from $B$ (and vice versa). Since $A\cap B=\emptyset$, we conclude $A^{*}_{\infty}\cap B^{*}_{\infty}=\emptyset$ and $|(A\cup B)^{*}_{\infty}|=|A^{*}_{\infty}|+|B^{*}_{\infty}|$. That means $\mu_{\infty}(A\cup B|Q)=\mu_{\infty}(A|Q)+\mu_{\infty}(B|Q)$ for computationally independent sets. That shows the proposed formula satisfies all required properties of semantic information measure.
All that brings us to the following conclusion. With infinite computing and reasoning power, total information encoded in $A$ must be preserved. Indeed, that bold (philosophical) claim is at the center of our investigations. Let us then justify it rigorously.
@@
We say that $X\subseteq\mathcal{K}$ and $Y\subseteq\mathcal{K}$ are equivalent in context $Q\subseteq\mathcal{K}$, i.e., $X\equiv Y$, if and only if $X\subseteq\mathcal{E}_{\infty}(Y|Q)$ and $Y\subseteq\mathcal{E}_{\infty}(X|Q)$.
Notice that the equivalence of knowledge does not mean $A$ is a fixed point, i.e., it is not true that $A=\mathcal{E}_{\infty}(A|Q)$ when $A\equiv\mathcal{E}_{\infty}(A|Q)$. If we have $A=\mathcal{E}_{\infty}(A|Q)$, then nothing more can be derived from $A$ in the context of $Q$ by any reasoning rule from $A$. That statement is very strong and is not true when $A\equiv\mathcal{E}_{\infty}(A|Q)$. Suppose $A=\{x\}$ and $f\in\mathcal{F}_{\infty}$ allows for getting $y=f(x)\in\mathcal{R}_{C}(Q)$. Then $\mathcal{E}_{\infty}(A|Q)=\{x,y\}\supset\{x\}=A\equiv\mathcal{E}_{\infty}(A|Q)$.
To proceed we will also need the idempotence of reasoning with unlimited computational budget.
Theorem 3 (Idempotence of unlimited reasoning).
For any $A\subseteq\mathcal{K}$ and context $Q\subseteq\mathcal{K}$ we have:
$\mathcal{E}_{\infty}(A|Q)=\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q)|Q).$ (12)
@@
By definition $\mathcal{E}_{\infty}(A|Q)\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))$. To show $\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))\subseteq\mathcal{E}_{\infty}(A|Q)$, we must find $h\in\mathcal{F}_{\infty}$ such that $h(A|Q)=g(A\cup\bigcup f_{i}(A|Q))$ for $g,f_{i}\in\mathcal{F}_{\infty}$. But since we can always find such $h\in\mathcal{F}_{\infty}$, we conclude that any context-relevant conclusion $g(\mathcal{E}_{\infty}(A|Q))$ derived from $\mathcal{E}_{\infty}(A|Q)$ by a reasoning rule $g\in\mathcal{F}_{\infty}$ can be also derived directly from $A$ as $h(A)=g(A\cup\bigcup f_{i}(A))$. Therefore, by definition of emergence operator we have $\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))\subseteq\mathcal{E}_{\infty}(A|Q)$ as required. ∎
With all the building blocks available, we can now to prove that unlimited reasoning ($C=\infty$) can reveal but never create information. The semantic content of a knowledge base is invariant under unlimited reasoning of emergence operator. Apparent information gain at finite computational budget emerges as an effect of limited computational access.
Theorem 4 (Information preservation).
For any $A\subseteq\mathcal{K}$ and context $Q\subseteq\mathcal{K}$, if the semantic information is constant on equivalent sets of knowledge, i.e., $\mu_{\infty}(X|Q)=\mu_{\infty}(Y|Q)$ whenever $X\equiv Y$, then:
$\mu_{\infty}(\mathcal{E}_{\infty}(A|Q)|Q)=\mu_{\infty}(A|Q).$ (13)
Proof.
We must show equivalence of the knowledge encoded in $A$ and $\mathcal{E}_{\infty}(A|Q)$, i.e., we demand $A\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))$. But, by Theorem 3, we see that $A\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))=\mathcal{E}_{\infty}(A|Q)$ and $\mathcal{E}_{\infty}(A|Q)\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))=\mathcal{E}_{\infty}(A|Q)$ as requested. Since $A\equiv\mathcal{E}_{\infty}(A|Q)$ and $\mu_{\infty}$ is constant on equivalent sets of knowledge by assumption, we conclude that $\mathcal{E}_{\infty}(A|Q)=\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))$. ∎
Justification for the Axiom of Unlimited Reasoning with Emergence Operator
The completeness (or closure property) of the set of reasoning functions $\mathcal{F}_{\infty}$ sounds somewhat mysterious, but is rather natural. We demand that for any reasoning steps $f$ and $g$, there exists a single step $h$ such that $g(A\cup f(A))=h(A)$. That is the nature of an idealized, complete reasoning system we want to investigate to understand its tendency to generate hallucinations.
The construct $\mathcal{F}_{\infty}$ is designed to represent the theoretical limit of reasoning with an unbounded computational capacity. It is not intended to model the practical, heuristic, and bounded reasoning of a real LLM, but rather the theoretical limit of what is ultimately knowable from a set of premises. Completeness is fundamental to formal systems, such as first-order logic or Turing-complete computations and Direct Revelation Principle in auction theory kuratowski_set_1978 ; sipser1996introduction ; fudenberg1991game . If one can prove a lemma $f(A)$ from a set of axioms $A$, and then prove a theorem $g(A\cup f(A))$ using that lemma, a direct proof of the theorem from the axioms alone $h(A)$ must exist by simply concatenating the proof steps. The composition of two computable functions is itself a computable function. A direct revelation mechanism in (auction) game theory is a composition of preference processing mechanisms where agents report their preferences directly to the mechanism, and the mechanism implements their preference processing strategies to generate outcome.
That axiom is essential for the proof of idempotence of unlimited reasoning in Theorem 3 Theorem 4. These are critical philosophical and mathematical results necessary for our study. With them we argue that the information gain observed in practical, constrained inference is an artifact of computational bounds, not the magical creation of knowledge ex nihilo. We acknowledge that depends on the assertion that any sequential reasoning process can be compiled into a single computational step. Therefore, the philosophical claim that information gain is an artifact of computational bounds rests on the acceptance of this structure of reasoning.

### v7

@@ 4.1.4 Limits of Knowledge and Emergence Operator
For any subset of knowledge $A\in\mathcal{U}(\mathcal{K})$ and context $Q\in\mathcal{B}(\mathcal{K})$ there exists $\mu_{\infty}(A|Q)=\lim_{C\to\infty}\mu_{C}(A|Q)$.
@@
Assume that $(A,Q)$ is fixed. Since the mapping $C\mapsto\mu_{C}(A|Q)$ is non-decreasing and bounded, the sequence $\{\mu_{C}(A|Q)\}_{C}$ converges with $C\to\infty$ by the monotone convergence theorem for real numbers. ∎
It may be surprising to discover that the existence of that limit does not prohibit emergent information. We can learn and obtain useful information without violating that fundamental limitation above. To explain that critical and rather controversial property of information processing, we first introduce the emergence operator $\mathcal{E}_{C}$ to describe knowledge becoming accessible through (computationally) bounded reasoning. Then we can study the implications in more details.
@@
The emergence operator is a mapping:
$\mathcal{E}_{C}(\cdot|Q)\colon\mathcal{A}(\mathcal{K})\mapsto\mathcal{A}(\mathcal{K})$ (20)
defined for a fixed $C>0$ and a Borel context $Q\in\mathcal{B}(\mathcal{K})$ as follows:
$\mathcal{E}_{C}(A|Q)=A\cup\bigcup_{f\in\mathcal{F}_{C}}\Big(f(A)\cap\mathcal{R}_{C}(Q)\Big).$ (21)
With infinite computational budget:
$\mathcal{E}_{\infty}(A|Q)=A\cup\bigcup_{f\in\mathcal{F}_{\infty}}\Big(f(A)\cap\mathcal{R}_{\infty}(Q)\Big).$ (22)
For any $n\geq 0$ and $C>0$ we define reasoning as a composition:
@@
Philosophically, the emergence operator describes reasoning as explicitation, i.e., as the process of making latent implications manifest. When $A\subseteq\mathcal{E}_{C}(A|Q)$, the discovered elements represent knowledge that was implicit in $A$ but required computational work $C$ to become accessible. That aligns with the pragmatist or constructivist view that understanding involves active construction. As we will see next, with new knowledge becoming accessible, new insights may become reachable, emerging from what was invisible in distributed subsets of initial knowledge.
Suppose $\mathcal{E}_{C}(A|Q)=A\cup B$. Then, we say that $B$ emerges from reasoning based on $A$, sets $A$ and $B$ are computationally dependent and we write $A\not\perp_{C}B$. In such a case, it follows from the properties of $\mu_{C}$ that we can gain useful information. Whenever $A\subseteq A\cup B$, we have:
@@
The emergence operator is monotonic, which means the reasoning process it describes is also monotonic:
$\text{if }A\subseteq B,\text{ then }\mathcal{E}_{C}(A|Q)\subseteq\mathcal{E}_{C}(B|Q).$ (25)
Also, we preserve insight monotonicity, because by definition with $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ for $C<D$, we have $\mathcal{E}_{C}(A|Q)\subseteq\mathcal{E}_{D}(A|Q)$. By design, it also guarantees existence of a fixed point to which the composition of reasoning rules converges.
The following theorem summarizes properties of the emergence operator.
Theorem 2 (Properties of the Emergence Operator).
The operator $\mathcal{E}_{C}(\cdot\mid Q)$ satisfies the following properties:
(closure) $\mathcal{E}_{C}(A|Q)\in\mathcal{A}(\mathcal{K})\text{ for all }A\in\mathcal{A}(\mathcal{K}),$ (26)
(inflation) $A\subseteq\mathcal{E}_{C}(A|Q),$ (27)
(monotonicity) $\mathcal{E}_{C}(A|Q)\subseteq\mathcal{E}_{C}(B|Q)\text{ if }A\subseteq B,$ (28)
(chain-continuous) $\mathcal{E}_{C}\Big(\bigcup_{n}A_{n}\Big|Q\Big)=\bigcup_{n}\mathcal{E}_{C}(A_{n}|Q)\text{ for any increasing chain }(A_{n})_{n\in\mathbb{N}}\subseteq\mathcal{A}(\mathcal{K}).$ (29)
Proof.
Since $f(A)$ is analytic (Borel measurable image of analytic set), $f(A)\cap\mathcal{R}_{C}(Q)$ is analytic, the union over the countable set $\mathcal{F}_{C}$ is analytic (countable union of analytic sets), we conclude $\mathcal{E}_{C}(A|Q)\in\mathcal{A}(\mathcal{K})$. Inflationary and monotone properties are immediate from the definition. Finally, the chain-continuity (or $\omega$-Scott-continuity) holds for $A_{1}\subseteq A_{2}\subseteq\dots\subseteq A_{n}$ because $f(\bigcup_{n}A_{n})=\bigcup_{n}f(A_{n})$ for any image generating function $f$. Namely, we have:
$\mathcal{E}_{C}\left(\bigcup_{n}A_{n}\Big|Q\right)$ $=\left(\bigcup_{n}A_{n}\right)\cup\bigcup_{f\in\mathcal{F}_{C}}\left(f\left(\bigcup_{n}A_{n}\right)\cap\mathcal{R}_{C}(Q)\right)$
$=\left(\bigcup_{n}A_{n}\right)\cup\bigcup_{f\in\mathcal{F}_{C}}\left(\left(\bigcup_{n}f(A_{n})\right)\cap\mathcal{R}_{C}(Q)\right)$
$=\left(\bigcup_{n}A_{n}\right)\cup\bigcup_{f\in\mathcal{F}_{C}}\bigcup_{n}\left(f(A_{n})\cap\mathcal{R}_{C}(Q)\right)$
$=\bigcup_{n\in\mathbb{N}}\left[A_{n}\cup\bigcup_{f\in\mathcal{F}_{C}}\left(f(A_{n})\cap\mathcal{R}_{C}(Q)\right)\right]$
$=\bigcup_{n\in\mathbb{N}}\mathcal{E}_{C}(A_{n}|Q).$
If we start reasoning with some knowledge, the reasoning process will yield at least as much (or more) knowledge. Clearly, that means we are considering settings in which reasoning does not invalidate previous conclusions when new premises are added. Also, applying the emergence operator to the infinite union yields the same result as applying the operator to each element of the union. That last property of chain-continuity is crucial for meaningful reasoning capable of reaching any consistent conclusion at all. Without it, as we show in Theorem 3, there is no guarantee that the iterative application of emergence operator will converge to consensus or a semantically equivalent (stable) response.
@@
For any $A=A_{0}\in\mathcal{A}(\mathcal{K})$ define $A_{n+1}=\mathcal{E}_{C}(A_{n}|Q)$ and consider the following closure operation:
$\mathrm{cl}_{\infty}(A|Q)=\bigcup_{n=0}^{\infty}A_{n}=\bigcup_{n=0}^{\infty}\mathcal{E}_{C}^{(n)}(A|Q).$ (30)
That closure defines a fixed point $A^{*}_{C}(Q)$ of the emergence operator, i.e.:
$A_{C}^{\ast}(Q)=\mathcal{E}_{C}(A_{C}^{\ast}|Q)\in\mathcal{A}(\mathcal{K}).$ (31)
With unlimited computational budget we have:
$A_{\infty}^{\ast}(Q)=\mathcal{E}_{\infty}(A_{\infty}^{\ast}|Q)\in\mathcal{A}(\mathcal{K}).$ (32)
@@
By the assumptions we have made, the reasoning takes place in the environment with the structure necessary for convergence. In the partially ordered set $(\mathcal{A}(\mathcal{K}),\subseteq)$ there exists a supremum of every countable chain of reasoning $A_{0}\subseteq A_{1}\subseteq A_{2}\subseteq\dots$ accumulating knowledge in analytic sets. That supremum is the countable union $\bigcup_{n=0}^{\infty}A_{n}=\mathrm{cl}_{\infty}(A|Q)$, which is also analytic set.
By Theorem 2, the emergence operator $\mathcal{E}_{C}$ is monotone and continuous on this partial order. Therefore, by the Tarski-Kleene fixed-point theorem Kleene1938 ; Tarski1955 the supremum $A_{C}^{\ast}$ is the least fixed point of $\mathcal{E}_{C}$ containing $A=A_{0}$, which holds for $C=\infty$ as well. ∎
If we start with some initial knowledge and repeatedly apply the reasoning with emergence operator $\mathcal{E}_{C}$, then the process of accumulating knowledge (potentially changing perspective and understanding) will eventually stabilize at a fixed point $A_{C}^{*}$. That fixed point represents everything that can be derived from initial knowledge with available reasoning power, and nothing more:
$A^{*}_{\infty}(Q)=A\cup\mathcal{E}_{\infty}(A|Q)\cup\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q)|Q)\cup\dots\cup\mathcal{E}^{(\infty)}_{\infty}(A|Q)=\mathrm{cl}_{\infty}(A|Q).$ (33)
We can establish an important property of that deductive closure operation.
Theorem 4 (Idempotence of Unlimited Reasoning).
The map $A\mapsto\mathrm{cl}_{\infty}(A|Q)$ is inflationary and idempotent on $\mathcal{A}(\mathcal{K})$, i.e.:
$A\subseteq\mathrm{cl}_{\infty}(A|Q)\quad\text{and}\quad\mathrm{cl}_{\infty}(\mathrm{cl}_{\infty}(A|Q)|Q)=\mathrm{cl}_{\infty}(A|Q).$ (34)
Proof.
By definition, the closure $\mathrm{cl}_{\infty}(A|Q)$ is the union of an iterative process that starts with $A_{0}=A$ and accumulates knowledge in $\bigcup_{n=0}^{\infty}A_{n}$. Therefore, $A\subseteq\mathrm{cl}_{\infty}(A\mid Q)$ and the map is inflationary.
Idempotence of the closure operation follows from Theorem 3. By the fixed-point property, we have:
$\mathrm{cl}_{\infty}(\mathrm{cl}_{\infty}(A|Q)|Q)$ $=\mathrm{cl}_{\infty}(A|Q)\cup\mathcal{E}_{\infty}(\mathrm{cl}_{\infty}(A|Q)|Q)\cup\dots\cup\mathcal{E}^{(\infty)}_{\infty}(\mathrm{cl}_{\infty}(A|Q)|Q)$ (35)
$=\mathrm{cl}_{\infty}(A|Q)\cup\mathrm{cl}_{\infty}(A|Q)\cup\dots\cup\mathrm{cl}_{\infty}(A|Q)=\mathrm{cl}_{\infty}(A|Q).$ (36)
Theorem 3 allows for introducing the following equivalence relation on the set of knowledge.
@@
We say that two subsets of knowledge, $A\in\mathcal{A}(\mathcal{K})$ and $B\in\mathcal{A}(\mathcal{K})$, are equivalent in the context $Q$, and denote that context equivalence as $A\equiv_{Q}B$, if and only if they have the same closure:
$A^{*}_{\infty}(Q)=B^{*}_{\infty}(Q).$ (37)
Equivalence of knowledge means that, given the same context $Q$, we reach the same conclusions (the least fixed points) starting from different states of knowledge, $A$ or $B$. Notice that any subset $A$ is equivalent to itself.
Notice that it is not true $A=\mathcal{E}_{\infty}(A|Q)$, when $A\equiv_{Q}\mathcal{E}_{\infty}(A|Q)$.
Example.
Consider $A=\{x\}$ and $f\in\mathcal{F}_{\infty}$ getting $y=f(x)\in\mathcal{R}_{\infty}(Q)$. Then, we see that $\mathcal{E}_{\infty}(A|Q)=\{x,y\}\supset\{x\}=A$ and $A\equiv\mathcal{E}_{\infty}(A|Q)$. Therefore, it is not true that $A=\mathcal{E}_{\infty}(A|Q)$ when $A\equiv\mathcal{E}_{\infty}(A|Q)$. ∎
The fixed-points of reasoning and the computationally unlimited emergence operator are the building blocks of the semantic information measure $\mu_{C}$. The following example shows how we can use $\mathcal{E}_{\infty}$ to construct $\mu_{\infty}$ that evaluates the contextual informational content of a given knowledge subset with unlimited insight of $C=\infty$.
Example.
Define:
$\mu_{\infty}(A|Q)=\log|\mathcal{K}|\cdot|A^{*}_{\infty}(Q)\cap\mathcal{R}_{\infty}(Q)|$
where $A^{*}_{\infty}(Q)=\bigcup_{n=0}^{|\mathcal{K}|}\mathcal{E}^{(n)}_{\infty}(A|Q)$ is the fixed point of reasoning. To improve readability, we write $A^{*}_{\infty}=A^{*}_{\infty}(Q)$.
The measure construction satisfies the required axioms of Definition 6. For the null set, we have $\mu_{\infty}(\emptyset|Q)=\log|\mathcal{K}|\cdot|\emptyset\cap\mathcal{R}_{\infty}(Q)|=0$. If $A\subseteq B$, then $A^{*}_{\infty}\subseteq B^{*}_{\infty}$ by monotonicity of $\mathcal{E}_{\infty}$, implying $\mu_{\infty}(A|Q)\leq\mu_{\infty}(B|Q)$. Contextual non-negativity holds since $\mu_{\infty}(A|Q)>0$ if and only if $A^{*}_{\infty}\cap\mathcal{R}_{\infty}(Q)\neq\emptyset$, which occurs precisely when $A$ contains or can derive relevant knowledge. The measure is bounded since $|A^{*}_{\infty}\cap\mathcal{R}_{\infty}(Q)|\leq|\mathcal{K}|$. For subadditivity, we have:
$\mu_{\infty}(A\cup B|Q)\leq\mu_{\infty}(A|Q)+\mu_{\infty}(B|Q),$
since $|(A\cup B)^{*}_{\infty}\cap\mathcal{R}_{\infty}(Q)|\leq|A^{*}_{\infty}\cap\mathcal{R}_{\infty}(Q)|+|B^{*}_{\infty}\cap\mathcal{R}_{\infty}(Q)|$.
Note that additivity of $\mu_{\infty}$ (equality) does not generally hold for computationally independent sets $A\perp_{\infty}B$. Even when neither $A$ nor $B$ can derive elements of the other, both may independently derive the same conclusions. For example, if $A=\{P,P\to R\}$ and $B=\{Q,Q\to R\}$, then $A\perp_{\infty}B$ yet both derive $R$, making $(A\cup B)^{*}_{\infty}\cap\mathcal{R}_{\infty}(Q)$ smaller than the sum of the individual closures. Thus, computational independence alone does not guarantee additivity of the semantic information measure. ∎
In that setting we can now ask the fundamental question:
What knowledge can we discover with unlimited computational budget?
The axioms we have introduced lead us to an interesting and intriguing conclusion. With infinite computing and reasoning power, total information must be preserved. That bold (philosophical) claim is at the center of our investigations. Let us prove it formally.
Theorem 5 (Information Preservation).
Assume that $\mu_{\infty}$ is invariant on equivalence classes, i.e., $A\equiv_{Q}B$ implies $\mu_{\infty}(A|Q)=\mu_{\infty}(B|Q)$. Then for all $A\in\mathcal{A}(\mathcal{K})$ we have:
$\mu_{\infty}(\mathrm{cl}_{\infty}(A|Q)|Q)=\mu_{\infty}(A|Q).$ (38)
@@
Since $A$ and $\mathrm{cl}_{\infty}(A|Q)$ have the same closure, they are equivalent by definition, $A\equiv_{Q}\mathrm{cl}_{\infty}(A\mid Q)$). Preservation of information follows from from the assumed invariance of $\mu_{\infty}$ on equivalence classes. ∎
Therefore, the answer to the question above is this. Unlimited reasoning does not create information, it can only reveal it and make accessible. All that can be known is already in $A$, we can only extract that knowledge and make useful.
That implies the apparent information gain at finite computational budget, $C<\infty$, emerges as an effect of limited computational access. Can we call that creativity? It is any increase in accessible information, as measured by $\mu_{C}$, under bounded computation.